OpenStudy (df001):

How is this answer obtained? 2pie*(3.6)/360=S/130 S=2.6pie

OpenStudy (df001):



OpenStudy (doc.brown):

It's pi

OpenStudy (doc.brown):

We know that there are 3.14 radians along the edge of half a circle, so to get the number of radians around the whole circle you have to multiply by 2.

OpenStudy (doc.brown):

The radius is half the diameter so one radian will have the length 7.2/2 = 3.6

OpenStudy (df001):

\[\frac{ 2\pi*3.6 }{ 360 }=\frac{ S }{ 130 }\]

OpenStudy (df001):

I do not get why the answer in pi form is 2.6

OpenStudy (df001):

Cross multiply fraction, 2940.530724/360 ; S=8.168140899

OpenStudy (df001):

In converted form of 8.168140899, shouldn't the answer be \[\frac{ 13\pi }{ 5 }\]

OpenStudy (doc.brown):

What you have here is a ratio. On the top you have the length of the circumference: two halves of a circle, with 3.14159 lengths of the radius in each half, and a radius of length 3.6. On the bottom you have another way of expressing the same thing. A circumference expressed in degrees, the circle divided into 360. They are equal to each other and can be divided to give you the number 1, like: \(\frac{2}{2}=1\) or \(\frac{beans}{beans}=1\)

OpenStudy (df001):

I get the radius is 3.6 and I know all the steps to solving this but, I can't seem to get S=2.6 on both my calculators

OpenStudy (doc.brown):

\(\frac{S}{130}\) is the same ratio and can be expressed as being equal to 1, or equal to \(\frac{2\pi(3.6)}{360}\). THe question being asked is how long is the length along the outside of the circle when the angle is equal to 130. Which of course is 180 - 50 shown in the picture.

OpenStudy (doc.brown):